# Smart Contracts in the Real World: A Statistical Exploration of External Data Dependencies

### v3

@@ 1 Introduction
The inherent decentralization of blockchain facilitates the execution of smart contracts with trust, eliminating the need for reliance on any single entity(Zhang et al., 2022). This pivotal characteristic has substantially driven the broad adoption of blockchain technology, notably within the Ethereum ecosystem. Consequently, this ecosystem has spurred the creation of a multitude of decentralized applications (DApps) to meet the escalating demand for their functionality(Li et al., 2020). The significance of external data for smart contracts is self-evident, given their frequent reliance on real-world data to enhance functionality. Engagement with real-time data, including prices or weather conditions, is commonplace for executing specific logic within smart contracts.
In this study, we examine 9,356 successfully compiled real-world smart contracts from an initial pool of 10,500, collected from Etherscan via a web crawler. We have excluded contracts that are outdated or contain syntax errors. Specifically, upon reviewing 240 smart contract source codes that interact with external data, we have compiled a keyword list titled ”oracle_services,” encompassing pertinent contract and function names. Furthermore, employing Abstract Syntax Trees (ASTs)(Huang et al., 2024), a tree-like data structure extensively utilized in compilers and interpreters to represent the abstract syntax structure of programming language code. We then compared the 9,356 contracts against the ”oracle_services” list, determining the number and proportion of contracts that involve external data interaction.
@@ 2.1 Blockchain and Ethereum
Blockchain(Guo and Yu, 2022), as a decentralized and distributed ledger, securely stores data across a network of computers, safeguarding against unauthorized tampering and modification. The security of blockchain has facilitated its widespread adoption across finance, healthcare, and various other sectors. Ethereum, a blockchain-based platform, empowers developers to craft and implement smart contracts and DApps. Established by Vitalik Buterin and his team in 2015, Ethereum introduced the programmable blockchain, enabling the storage and execution of both transactions and executable code. Ethereum’s flexible, Turing-complete programming environment positions it as developers’ top choice for constructing decentralized solutions.
@@ 2.2 Smart Contracts and Audit Reports
Smart contracts(Tolmach et al., 2021) are Turing-complete programs executed on blockchain infrastructure. They are commonly implemented using languages such as Solidity, providing functionalities applicable across diverse real-world domains, including finance, cryptography, and healthcare(Perez and Livshits, 2021). Their widespread deployment in financial applications renders them a frequent target for hacking endeavors. Owing to their immutable character post-deployment, security firms provide audit services to mitigate associated risks. Smart contract security audits(Li et al., 2021) encompass manual analyses conducted by specialized security teams to identify and assess potential vulnerabilities. Following their execution on simulated blockchains, audit reports are disseminated, encompassing details such as contract addresses, user permissions, specific vulnerabilities, and corrective recommendations(Li et al., 2023b). These reports are conventionally presented in English to accommodate a global readership.
@@ 2.3 External Data Dependencies
The external data dependency(Torres et al., 2021) of smart contracts pertains to scenarios where a smart contract necessitates interaction with data sources external to the blockchain during execution. Typically, smart contracts necessitate the retrieval of data from external sources. This retrieval of external data is essential for smart contract execution, enabling them to make decisions predicated on real-time or dynamic external conditions. Contracts engage with off-chain data via Oracle services. Oracle services(Caldarelli, 2022) for smart contracts function as intermediaries, supplying external data to blockchain-based smart contracts and facilitating communication with real-world data. Furthermore, cross-chain communication(Shao et al., 2021) is a prevalent method for smart contracts to interact with external data, frequently employing technologies such as cross-chain bridges, notarization mechanisms, relay chain protocols, and others. In this article, we present several real-world instances where smart contract interactions with external data have precipitated security concerns.
@@ 3.4 Analysis
Concurrently with our ASTs analysis, we acquired and analyzed 249 security audit reports pertaining to smart contracts. These reports offer valuable insights into the frequency of external data access across diverse projects. Our statistical analysis yields a spectrum of external data access frequencies, with the maximum number of accesses recorded at an astonishing 11,724 instances. The average frequency of external data access among the audited projects was calculated to be 117 occurrences per project. To elucidate the patterns of external data access within smart contracts, we utilized a visualization technique to depict the distribution of access frequencies graphically. Based on the observed distribution, we applied a decision tree classification method to establish two critical thresholds, facilitating the categorization of access frequencies into three distinct levels: low, medium, and high. This stratification approach not only offers a clear overview of project distribution but also enhances the understanding of the dependency on and variability of external data within the smart contract ecosystem.
@@
Subsequently, we compiled and analyzed 74 smart contracts from a total of 249 audit reports, transforming them into ASTs and subsequently into CFG. Applying the cyclomatic complexity formula: $V(G)=E-N+2$ (where $E$ represents the edges and $N$ represents the nodes in the CFG), the process is detailed in the accompanying Algorithm 1. This methodology addresses the challenge posed by the lack of existing libraries for computing the cyclomatic complexity of Solidity contracts. We observed a significant positive correlation (with a correlation coefficient $r$ and a $p-value<0.5$) between the number of external data dependencies and the code complexity of smart contracts, as depicted in Fig.4
@@
Within the range of low dependency counts, the increase in code complexity is relatively gradual; however, with an increasing number of dependencies, the code complexity exhibits a trend of accelerating growth. Several factors may contribute to the observed positive correlation(Jahncke, 2023). Firstly, an increased dependence on external data may require incorporating more complex logical judgments and state management mechanisms within the contract. Secondly, to process and verify external data, the contract may need to include a more significant number of conditional statements and loop structures, thereby increasing the cyclomatic complexity of the code.
@@ 4.1.1 DeFi
DeFi(Busayatananphon and Boonchieng, 2022) is among the most prominent applications of smart contract technology. Decentralized oracles are predominant in this sector, supplying dependable external data for decentralized exchanges, lending platforms, and other financial instruments. Moreover, sidechain and cross-chain technologies are crucial for lowering transaction costs and enabling asset interoperability among different blockchain networks. With escalating privacy concerns, ZKPs technology is increasingly adopted in the DeFi sector, particularly for scenarios that necessitate safeguarding transaction details and account balances.
@@ 4.1.2 Supply Chain Management
In supply chain management(Chandan et al., 2023), smart contracts are mainly used to optimize logistics and inventory systems. High transaction and data processing demands have led to the widespread adoption of sidechain technology, which effectively reduces the load on the main blockchain and enhances system efficiency. Furthermore, smart contracts in this domain heavily depend on Oracle technology to obtain external data like market prices and shipping statuses, ensuring accuracy and real-time responsiveness. Though the use of ZKPs is still limited, their capacity to bolster data privacy is gaining traction, especially in scenarios with sensitive commercial data.
@@ 4.1.3 Gaming
The gaming industry has a significant need for external data interaction(Jiang et al., 2022), particularly for event triggers, random number generation, and verifying player behavior. Centralized oracles are essential in this domain, providing an efficient and simple method for accessing external data. Sidechain technology is often employed in blockchain-based gaming ecosystems to alleviate the main blockchain’s load and reduce transaction costs. Despite their limited prevalence in the gaming sector, cross-chain technology and ZKPs show promise in certain scenarios, including facilitating interoperability across various blockchain networks and safeguarding sensitive gaming data.
@@ 4.2.1 Centralized Oracle
Centralized oracles(Pasdar et al., 2023), as singular entities, supply external data to blockchain networks and act as the authoritative sources for information like market prices and real-world events. Despite their efficiency and simplicity, centralized oracles present a single point of failure and potential trust issues because their data feeds can be manipulated or experience outages. Relying on a single data source can compromise the decentralized nature of blockchain systems.
The centralized oracle serves as an intermediary within the blockchain system, tasked with integrating external data into on-chain smart contracts. The typical implementation process encompasses the following stages(Cai et al., 2022): Initially, the oracle retrieves necessary information from external sources such as APIs, databases, or sensors, ensuring the data’s completeness and timeliness during collection. Next, the oracle validates the collected data for accuracy and completeness, serving as a reliable data source for executing smart contracts. Following this, the verified data is transmitted to the blockchain network via the oracle server, thereby finalizing the on-chain data transmission and dissemination. Eventually, the smart contract receives the data provided by the prediction machine and executes the corresponding operation according to the prediction logic, thus realizing the efficient interaction between external information and the blockchain system.
The sample code is illustrated in Listing 1. Within this contract, the oracle address acts as the centralized predicate machine, and it alone can execute the ’updateData’ function to modify the data on-chain. However, centralized oracles are prone to a single point of failure and trust issues, potentially compromising the decentralized nature and security of blockchain systems(Košt’ál et al., 2023). Thus, despite the simplicity of implementing centralized oracles, their associated risks must be balanced against system requirements in practical applications.
@@ 4.2.2 Decentralized Oracle
Decentralized oracles(Mustafa et al., 2023) reduce the risks of centralization by consolidating data from various sources, which increases reliability and trustworthiness. They use consensus mechanisms to authenticate information prior to its transmission to smart contracts, minimizing the chance of data manipulation. This approach is in line with blockchain’s foundational principles, ensuring that no single entity exerts excessive control over the data.
During implementation, a smart contract initially requests specific data, received by several independent oracle nodes. In response to the request, each node gathers the necessary information from its designated external data sources and then submits this data to the blockchain. Subsequently, the data submitted by all the nodes are verified and aggregated through a predetermined consensus mechanism (e.g., majority voting, weighted average, etc.) to ensure the accuracy and integrity of the data. Ultimately, the consolidated data results are transmitted to the smart contract, triggering the execution of its logic and facilitating the secure interaction between the blockchain system and external data(Xiong et al., 2024).
The sample code is illustrated in Listing 2. Within this contract, the ’AggregatorV3Interface’ interface facilitates interaction with decentralized oracles, including Chainlink’s data aggregator, to retrieve the most recent price data. Decentralized oracles mitigate trust risks and enhance data reliability and accuracy by incorporating multiple independent data sources and nodes(Gupta et al., 2024). As noted in the Chainlink whitepaper, “Existing oracles rely on a centralized service model, which can lead to a single point of failure in smart contracts, yet they are considered more secure than traditional centralized digital contracts.”
@@ 4.2.3 Sidechain Technology
Sidechains(Soltani et al., 2022) are separate blockchains that function alongside a mainchain, facilitating the transfer of assets and data between them. They promote scalability and flexibility by transferring transactions away from the mainchain, thereby reducing network congestion and improving performance. Sidechains can be tailored to specific use cases, offering a sandbox for testing new features without endangering the security of the main network. The typical implementation process encompasses the following steps(Yadav et al., 2022):
@@
On the sidechain, the necessary verification and minting logic must be implemented to ensure the secure transfer of assets. Notably, sidechain implementation must address the complexities of cross-chain communication, security, and consensus mechanisms. For instance, the Zendoo protocol offers a verifiable cross-chain transfer protocol utilizing zk-SNARKs, enabling the main chain to engage with various sidechains without requiring knowledge of their internal structures(Balani and Chavan, 2023). The protocol generates succinct proofs of sidechain state progression through the recursive combination of zk-SNARKs, allowing the main chain to efficiently validate sidechain transactions.
@@ 4.2.4 Cross-chain Technology
Cross-chain technology(Li et al., 2023a) enables interoperability among various blockchain networks, allowing for the frictionless transfer of assets and data across different platforms. Such interoperability is essential for fostering a unified blockchain ecosystem, enabling users to harness the unique advantages of different networks. Methods like atomic swaps, relays, and decentralized oracles are used for cross-chain communication, each presenting distinct trade-offs in terms of security, speed, and complexity.
Cross-chain technology is designed to facilitate interoperability among diverse blockchain networks, enabling the frictionless transfer and interaction of assets and data across various chains. The typical implementation process encompasses the following steps(Zhong et al., 2022):
@@
    Cross-chain communication protocols: Protocols are standardized to facilitate secure and reliable information exchange between disparate blockchain networks. They stipulate message formats, transmission mechanisms, and authentication procedures.
@@
    Atomicity of Cross-Chain Transactions: Ensure that cross-chain transactions are atomic, meaning they either execute successfully across all involved chains or are fully rolled back to avoid partial completions that could result in inconsistencies. This is commonly achieved using atomic swaps or complementary protocols.
@@
    Relay and Verification Mechanism: Implement relay nodes or verifiers tasked with monitoring various blockchains’ statuses and facilitating the transfer and verification of transaction data between chains, ensuring the security and precision of cross-chain operations.
@@
    Bridging: Establish a mechanism for cross-chain bridging that enables the transfer of assets between distinct blockchains. This process generally involves locking assets on the originating chain and issuing an equivalent token on the destination chain.
@@ 4.2.5 ZKPs (Zero-Knowledge Proof) Data Interaction
ZKPs(Lavin et al., 2024) are cryptographic techniques that enable a party to verify the truth of a statement without disclosing extra information. In blockchain applications, ZKPs bolster privacy and security by allowing verification of data without exposing the underlying data. This is especially advantageous in confidential scenarios like identity verification and private transactions, ensuring sensitive data stays safeguarded during necessary validation processes.
In implementing a ZKPs protocol, the prover and verifier initially agree on the protocol to be used and establish the required public parameters, ensuring the protocol’s framework and fairness. Within the protocol’s framework, the prover must define the statements to be proven and translate them into mathematical forms compatible with ZKPs for further processing. Subsequently, the prover creates a proof of a specific mathematical model or rule, a process that often requires intricate calculations to validate the statement’s truth without revealing sensitive data. The prover subsequently submits the proof to the verifier for validation. Finally, the verifier assesses the proofs using the protocol’s specified verification methods to ascertain the statement’s validity and ensure compliance with the ZKPs protocol’s requirements(Gabay et al., 2020).

### v4

@@ 1 Introduction
The inherent decentralization of blockchain facilitates the execution of smart contracts with trust, eliminating the need for reliance on any single entity [59]. This pivotal characteristic has substantially driven the broad adoption of blockchain technology, notably within the Ethereum ecosystem. Consequently, this ecosystem has spurred the creation of a multitude of decentralized applications (DApps) [42, 21, 44] to meet the escalating demand for their functionality [27]. The significance of external data for smart contracts is self-evident, given their frequent reliance on real-world data to enhance functionality. Engagement with real-time data, including prices or weather conditions, is commonplace for executing specific logic within smart contracts.
In this study, we examine 9,356 successfully compiled real-world smart contracts from an initial pool of 10,500, collected from Etherscan via a web crawler. We have excluded contracts that are outdated or contain syntax errors. Specifically, upon reviewing 240 smart contract source codes that interact with external data, we have compiled a keyword list titled ”oracle_services,” encompassing pertinent contract and function names. Furthermore, employing Abstract Syntax Trees (ASTs) [17, 38], a tree-like data structure extensively utilized in compilers and interpreters to represent the abstract syntax structure of programming language code. We then compared the 9,356 contracts against the ”oracle_services” list, determining the number and proportion of contracts that involve external data interaction.
@@ 2.1 Blockchain and Ethereum
Blockchain [14, 12, 52, 36], as a decentralized and distributed ledger, securely stores data across a network of computers, safeguarding against unauthorized tampering and modification. The security of blockchain has facilitated its widespread adoption across finance, healthcare, and various other sectors. Ethereum, a blockchain-based platform, empowers developers to craft and implement smart contracts and DApps. Established by Vitalik Buterin and his team in 2015, Ethereum introduced the programmable blockchain, enabling the storage and execution of both transactions and executable code. Ethereum’s flexible, Turing-complete programming environment positions it as developers’ top choice for constructing decentralized solutions.
@@ 2.2 Smart Contracts and Audit Reports
Smart contracts [48, 29, 51] are Turing-complete programs executed on blockchain infrastructure. They are commonly implemented using languages such as Solidity, providing functionalities applicable across diverse real-world domains, including finance, cryptography, and healthcare [41]. Their widespread deployment in financial applications renders them a frequent target for hacking endeavors. Owing to their immutable character post-deployment, security firms provide audit services to mitigate associated risks. Smart contract security audits [26, 53, 22] encompass manual analyses conducted by specialized security teams to identify and assess potential vulnerabilities. Following their execution on simulated blockchains, audit reports are disseminated, encompassing details such as contract addresses, user permissions, specific vulnerabilities, and corrective recommendations [28]. These reports are conventionally presented in English to accommodate a global readership.
@@ 2.3 External Data Dependencies
The external data dependency [49] of smart contracts pertains to scenarios where a smart contract necessitates interaction with data sources external to the blockchain during execution. Typically, smart contracts necessitate the retrieval of data from external sources. This retrieval of external data is essential for smart contract execution, enabling them to make decisions predicated on real-time or dynamic external conditions. Contracts engage with off-chain data via Oracle services. Oracle services [5, 8] for smart contracts function as intermediaries, supplying external data to blockchain-based smart contracts and facilitating communication with real-world data. Furthermore, cross-chain communication [45, 11, 7] is a prevalent method for smart contracts to interact with external data, frequently employing technologies such as cross-chain bridges, notarization mechanisms, relay chain protocols, and others. In this article, we present several real-world instances where smart contract interactions with external data have precipitated security concerns.
@@ 3.4 Analysis
Concurrently with our ASTs analysis, we acquired and analyzed 249 security audit reports pertaining to smart contracts. These reports offer valuable insights into the frequency of external data access across diverse projects. Our statistical analysis yields a spectrum of external data access frequencies, with the maximum number of accesses recorded at an astonishing 11,724 instances. The average frequency of external data access among the audited projects was calculated to be 117 occurrences per project. To elucidate the patterns of external data access within smart contracts, we utilized a visualization technique to depict the distribution of access frequencies graphically. Based on the observed distribution, we applied a decision tree classification method [32, 34, 1] to establish two critical thresholds, facilitating the categorization of access frequencies into three distinct levels: low, medium, and high. This stratification approach not only offers a clear overview of project distribution but also enhances the understanding of the dependency on and variability of external data within the smart contract ecosystem.
@@
Subsequently, we compiled and analyzed 74 smart contracts from a total of 249 audit reports, transforming them into ASTs and subsequently into CFG. Applying the cyclomatic complexity formula: V(G)=EN+2V(G)=E-N+2 (where $E$ represents the edges and $N$ represents the nodes in the CFG), the process is detailed in the accompanying Algorithm 1. This methodology addresses the challenge posed by the lack of existing libraries for computing the cyclomatic complexity of Solidity contracts. We observed a significant positive correlation (with a correlation coefficient $r$ and a pvalue<0.5p-value<0.5) between the number of external data dependencies and the code complexity of smart contracts, as depicted in Fig.4
@@
Within the range of low dependency counts, the increase in code complexity is relatively gradual; however, with an increasing number of dependencies, the code complexity exhibits a trend of accelerating growth. Several factors may contribute to the observed positive correlation[18]. Firstly, an increased dependence on external data may require incorporating more complex logical judgments and state management mechanisms within the contract. Secondly, to process and verify external data, the contract may need to include a more significant number of conditional statements and loop structures, thereby increasing the cyclomatic complexity of the code.
@@ 4.1.1 DeFi
DeFi[3] is among the most prominent applications of smart contract technology. Decentralized oracles are predominant in this sector, supplying dependable external data for decentralized exchanges, lending platforms, and other financial instruments. Moreover, sidechain and cross-chain technologies are crucial for lowering transaction costs and enabling asset interoperability among different blockchain networks. With escalating privacy concerns, ZKPs technology is increasingly adopted in the DeFi sector, particularly for scenarios that necessitate safeguarding transaction details and account balances.
@@ 4.1.2 Supply Chain Management
In supply chain management[6], smart contracts are mainly used to optimize logistics and inventory systems. High transaction and data processing demands have led to the widespread adoption of sidechain technology, which effectively reduces the load on the main blockchain and enhances system efficiency. Furthermore, smart contracts in this domain heavily depend on Oracle technology to obtain external data like market prices and shipping statuses, ensuring accuracy and real-time responsiveness. Though the use of ZKPs is still limited, their capacity to bolster data privacy is gaining traction, especially in scenarios with sensitive commercial data.
@@ 4.1.3 Gaming
The gaming industry has a significant need for external data interaction[19], particularly for event triggers, random number generation, and verifying player behavior. Centralized oracles are essential in this domain, providing an efficient and simple method for accessing external data. Sidechain technology is often employed in blockchain-based gaming ecosystems to alleviate the main blockchain’s load and reduce transaction costs. Despite their limited prevalence in the gaming sector, cross-chain technology and ZKPs show promise in certain scenarios, including facilitating interoperability across various blockchain networks and safeguarding sensitive gaming data.
@@ 4.2.1 Centralized Oracle
Centralized oracles[40], as singular entities, supply external data to blockchain networks and act as the authoritative sources for information like market prices and real-world events. Despite their efficiency and simplicity, centralized oracles present a single point of failure and potential trust issues because their data feeds can be manipulated or experience outages. Relying on a single data source can compromise the decentralized nature of blockchain systems.
The centralized oracle serves as an intermediary within the blockchain system, tasked with integrating external data into on-chain smart contracts. The typical implementation process encompasses the following stages[4]: Initially, the oracle retrieves necessary information from external sources such as APIs, databases, or sensors, ensuring the data’s completeness and timeliness during collection. Next, the oracle validates the collected data for accuracy and completeness, serving as a reliable data source for executing smart contracts. Following this, the verified data is transmitted to the blockchain network via the oracle server, thereby finalizing the on-chain data transmission and dissemination. Eventually, the smart contract receives the data provided by the prediction machine and executes the corresponding operation according to the prediction logic, thus realizing the efficient interaction between external information and the blockchain system.
The sample code is illustrated in Listing 1. Within this contract, the oracle address acts as the centralized predicate machine, and it alone can execute the ’updateData’ function to modify the data on-chain. However, centralized oracles are prone to a single point of failure and trust issues, potentially compromising the decentralized nature and security of blockchain systems[20]. Thus, despite the simplicity of implementing centralized oracles, their associated risks must be balanced against system requirements in practical applications.
@@ 4.2.2 Decentralized Oracle
Decentralized oracles[37] reduce the risks of centralization by consolidating data from various sources, which increases reliability and trustworthiness. They use consensus mechanisms to authenticate information prior to its transmission to smart contracts, minimizing the chance of data manipulation. This approach is in line with blockchain’s foundational principles, ensuring that no single entity exerts excessive control over the data.
During implementation, a smart contract initially requests specific data, received by several independent oracle nodes. In response to the request, each node gathers the necessary information from its designated external data sources and then submits this data to the blockchain. Subsequently, the data submitted by all the nodes are verified and aggregated through a predetermined consensus mechanism (e.g., majority voting, weighted average, etc.) to ensure the accuracy and integrity of the data. Ultimately, the consolidated data results are transmitted to the smart contract, triggering the execution of its logic and facilitating the secure interaction between the blockchain system and external data[55].
The sample code is illustrated in Listing 2. Within this contract, the ’AggregatorV3Interface’ interface facilitates interaction with decentralized oracles, including Chainlink’s data aggregator, to retrieve the most recent price data. Decentralized oracles mitigate trust risks and enhance data reliability and accuracy by incorporating multiple independent data sources and nodes[15]. As noted in the Chainlink whitepaper, “Existing oracles rely on a centralized service model, which can lead to a single point of failure in smart contracts, yet they are considered more secure than traditional centralized digital contracts.”
@@ 4.2.3 Sidechain Technology
Sidechains[46] are separate blockchains that function alongside a mainchain, facilitating the transfer of assets and data between them. They promote scalability and flexibility by transferring transactions away from the mainchain, thereby reducing network congestion and improving performance. Sidechains can be tailored to specific use cases, offering a sandbox for testing new features without endangering the security of the main network. The typical implementation process encompasses the following steps[57]:
@@
On the sidechain, the necessary verification and minting logic must be implemented to ensure the secure transfer of assets. Notably, sidechain implementation must address the complexities of cross-chain communication, security, and consensus mechanisms. For instance, the Zendoo protocol offers a verifiable cross-chain transfer protocol utilizing zk-SNARKs, enabling the main chain to engage with various sidechains without requiring knowledge of their internal structures[2]. The protocol generates succinct proofs of sidechain state progression through the recursive combination of zk-SNARKs, allowing the main chain to efficiently validate sidechain transactions.
@@ 4.2.4 Cross-chain Technology
Cross-chain technology[25] enables interoperability among various blockchain networks, allowing for the frictionless transfer of assets and data across different platforms. Such interoperability is essential for fostering a unified blockchain ecosystem, enabling users to harness the unique advantages of different networks. Methods like atomic swaps, relays, and decentralized oracles are used for cross-chain communication, each presenting distinct trade-offs in terms of security, speed, and complexity.
Cross-chain technology is designed to facilitate interoperability among diverse blockchain networks, enabling the frictionless transfer and interaction of assets and data across various chains. The typical implementation process encompasses the following steps[61]:
@@
    Cross-chain communication protocols: Protocols are standardized to facilitate secure and reliable information exchange between disparate blockchain networks [35] . They stipulate message formats, transmission mechanisms, and authentication procedures.
@@
    Atomicity of Cross-Chain Transactions: Ensure that cross-chain transactions are atomic, meaning they either execute successfully across all involved chains or are fully rolled back to avoid partial completions that could result in inconsistencies [9]. This is commonly achieved using atomic swaps or complementary protocols.
@@
    Relay and Verification Mechanism: Implement relay nodes or verifiers tasked with monitoring various blockchains’ statuses and facilitating the transfer and verification of transaction data between chains, ensuring the security and precision of cross-chain operations [56] .
@@
    Bridging: Establish a mechanism for cross-chain bridging that enables the transfer of assets between distinct blockchains [58] . This process generally involves locking assets on the originating chain and issuing an equivalent token on the destination chain.
@@ 4.2.5 ZKPs (Zero-Knowledge Proof) Data Interaction
ZKPs [23, 24, 54, 16] are cryptographic techniques that enable a party to verify the truth of a statement without disclosing extra information. In blockchain applications, ZKPs bolster privacy and security by allowing verification of data without exposing the underlying data. This is especially advantageous in confidential scenarios like identity verification and private transactions, ensuring sensitive data stays safeguarded during necessary validation processes.
In implementing a ZKPs protocol, the prover and verifier initially agree on the protocol to be used and establish the required public parameters, ensuring the protocol’s framework and fairness. Within the protocol’s framework, the prover must define the statements to be proven and translate them into mathematical forms compatible with ZKPs for further processing. Subsequently, the prover creates a proof of a specific mathematical model or rule, a process that often requires intricate calculations to validate the statement’s truth without revealing sensitive data. The prover subsequently submits the proof to the verifier for validation. Finally, the verifier assesses the proofs using the protocol’s specified verification methods to ascertain the statement’s validity and ensure compliance with the ZKPs protocol’s requirements[13].